OpenStudy (anonymous):

How many moles of dipyrithione (C10H8N2O2S2) contain 8.2 x 10^24 atoms of N2?

OpenStudy (anonymous):

Use the avogadro number, to divide your number of atoms N/Na = n (in mol) Na=6.022* 10^23 N= number of atoms n is amount of substance you're searching

OpenStudy (anonymous):

I divided 8.2x10^24 by .11(the percent mass of N2 in dipyrithione) to get the number if atoms in the entire compound then I divided by 6.02 x10^23 to get the moles of dipyrithione. I ended up with like 123. I the answer in my textbook says 6.8 though.

OpenStudy (anonymous):

I answered too fast i'm sorry, don't mind the percent mass, all you have to do is divide the number of atom by NA, which will give you the amount of substance of N, since you're looking for N2 you have to divide that by 2 which will give you the answer 6.81: 8.2x10^24 / 6.022* 10^23 = 13.6 13.6/2=6.8

OpenStudy (anonymous):

Thanks for the help!